# SAMA-UNet: Enhancing Medical Image Segmentation with Self-Adaptive Mamba-Like Attention and Causal-Resonance Learning

### v1

@@ 4.5 Parameters and computational efficiency
We also evaluate and compare the computational complexity of each model by measuring the number of trainable parameters and floating-point operations (FLOPs) using the BTCV dataset. As presented in Table 4, our proposed method, SAMA-UNet, achieves a balanced trade-off with 29.03M parameters and 145.67 GFLOPs, maintaining computational efficiency while delivering competitive performance. Compared to other state-of-the-art models, SAMA-UNet requires significantly fewer parameters and FLOPs than heavy transformer-based models such as LKM-UNet (189.55M, 993.70G) and SwinUNETR (100.44M, 281.69G). Even against models with encoder-only adaptations like U-Mamba (Enc) and nnUNet, our method remains the most efficient, indicating faster inference potential and reduced memory overhead. These results highlight the computational advantage of SAMA-UNet, making it a suitable choice for real-time or resource-constrained medical imaging applications without compromising segmentation quality.
@@ 5.1 Technical impact
In contrast to natural language, image tokens in medical segmentation tasks do not follow an inherently causal sequence. Traditional SSMs, specifically designed for autoregressive modelling, face challenges in effectively capturing spatial dependencies in this context. To address this, we replace SSM-based token mixers in the encoder with our proposed SAMA block. SAMA introduces a hybrid token mixing strategy that utilizes both pixel-focused softmax attention and mamba-inspired macro architectural elements that allow the model to capture both local details and global semantic dependencies without enforcing a causal sequence. Additionally, we integrate differential attention and flash attention mechanisms. The mechanisms inspired by the human foveal vision system allow the network to focus on the most relevant regions while maintaining linear computational complexity. Our experiments, shown in Figures 4 and 5, confirm that using the SAMA block instead of SSMs leads to significant performance gains while also reducing the number of model parameters and FLOPs compared to designs based on SSMs. While attention mechanisms help mix tokens in the encoder, U-shaped segmentation networks also depend a lot on skip connections to share features at different scales. Here, we introduced the CR-MSM to capture implicit causal relationships across resolution levels. In U-shaped networks, encoder features at higher resolutions inherently precede coarser representations and often provide foundational cues for decoder reconstruction. However, conventional multi-scale fusion (e.g., simple concatenation or convolution-based aggregation) fails to preserve this directional causality. CR-MSM addresses this concern by applying four directional transformations (original, transposed, flipped, and flipped-transposed) to the encoder features. A shared SSM block flattens and processes each transformed view, capturing long-range dependencies across spatial orientations. To maintain causal alignment during fusion, we introduce a causal resonance strategy that performs directional averaging rather than naive concatenation. This fusion ensures the model respects the natural order of feature generation.
In the ablation study, Table III confirms the importance of each component. Replacing causal fusion with concatenation results in a decrement in DSC; removing multi-view transformations causes a drop. Together, these results validate our hypothesis that spatial causality is crucial for preserving semantic continuity in medical images.
@@ 5.3 Limitations and future directions
However, several challenges remain. First, integrating the Mamba macro architecture into the SAMA module increases the parameter count that necessitates more effective parameter optimization strategies. Second, expanding the model to handle 3D image segmentation uses a lot more memory because it triples the number of scanning sequences in the Vision State Space module. Addressing these challenges requires designing modules that balance efficiency and effectiveness, particularly for 3D network applications. Future work will focus on refining the network architecture, optimizing parameter utilization, and exploring real-time segmentation in large-scale datasets.
@@ 6 Conclusions
In this work, we introduced a novel SAMA-UNet architecture for medical image segmentation that effectively combines the strengths of dynamic attention, Mamba-based modeling, and multi-scale feature fusion. Our design uses the Self-Adaptive Mamba-Like Aggregated Attention block to improve how well the model understands both global and local features while keeping the calculations efficient. Additionally, we proposed the Causal-Resonance Multi-Scale Module to facilitate causally aligned feature integration across encoder–decoder stages. Through extensive evaluations on four diverse benchmarks—BTCV, ACDC, EndoVis17, and ATLAS23—SAMA-UNet outperformed state-of-the-art CNN, Transformer, and Mamba-based models and achieved DSC scores of 85.38%, 92.16%, 67.14%, and 84.06%, and NSD scores of 87.82%, 96.54%, 68.70%, and 88.47%, respectively. These results validate the robustness and generalizability of our model across different modalities, including CT, MRI, and endoscopic imaging. Despite its strong performance, SAMA-UNet has limitations. The integration of Mamba structures is beneficial for sequence modeling but introduces additional parameters. Furthermore, extending the architecture to 3D segmentation scenarios leads to increased memory usage due to the enlarged SS2D scanning space. Addressing these challenges, particularly through lightweight design optimization and efficient 3D adaptations, will guide our future research efforts.
@@
Acknowledgements
This work was supported and funded by the Deanship of Scientific Research at Imam Mohammad Ibn Saud Islamic University (IMSIU) (grant number IMSIU-DDRSP2501)

### v2

@@ 4.5 Parameters and computational efficiency
We also evaluate and compare the computational complexity of each model by measuring the number of trainable parameters and floating-point operations (FLOPs) using the BTCV dataset. As presented in Table 4, our proposed SAMA-UNet achieves a balanced trade-off with 29.03M parameters and 145.67 GFLOPs, maintaining computational efficiency while delivering competitive performance. In contrast, removing the CR-MSM reduces the parameters to 27.2M but increases the computational cost to 356 GFLOPs. Compared to other state-of-the-art models, SAMA-UNet requires significantly fewer parameters and FLOPs than heavy transformer-based models such as LKM-UNet has 189.55M and 993.70G while SwinUNETR has 100.44M and 281.69G. Even against models with encoder-only adaptations like U-Mamba (Enc) and nnUNet, our method remains the most efficient, indicating faster inference potential and reduced memory overhead. These results highlight the computational advantage of SAMA-UNet, making it a suitable choice for real-time or resource-constrained medical imaging applications without compromising segmentation quality.
@@ 5.1 Technical impact
In contrast to natural language, image tokens in medical segmentation tasks do not follow an inherently causal sequence. Traditional SSMs, specifically designed for autoregressive modelling, face challenges in effectively capturing spatial dependencies in this context. To address this, we replace SSM-based token mixers in the encoder with our proposed SAMA block. SAMA introduces a hybrid token mixing strategy that utilizes both pixel-focused softmax attention and mamba-inspired macro architectural elements that allow the model to capture both local details and global semantic dependencies without enforcing a causal sequence. Additionally, we integrate differential attention and flash attention mechanisms. The mechanisms inspired by the human foveal vision system allow the network to focus on the most relevant regions while maintaining linear computational complexity. Our experiments, shown in Figures 4 and 5, confirm that using the SAMA block instead of SSMs leads to significant performance gains while also reducing the number of model parameters and FLOPs compared to designs based on SSMs. While attention mechanisms help mix tokens in the encoder, U-shaped segmentation networks also depend a lot on skip connections to share features at different scales. Here, we introduced the CR-MSM to capture implicit causal relationships across resolution levels. In U-shaped networks, encoder features at higher resolutions inherently precede coarser representations and often provide foundational cues for decoder reconstruction. However, conventional multi-scale fusion such as simple concatenation or convolution-based aggregation fails to preserve this directional causality. CR-MSM addresses this concern by applying four directional transformations to the encoder features. A shared SSM block flattens and processes each transformed view, capturing long-range dependencies across spatial orientations. To maintain causal alignment during fusion, we introduce a causal resonance strategy that performs directional averaging rather than naive concatenation. This fusion ensures the model respects the natural order of feature generation.
In the ablation study, Table 3 confirms the importance of each component. Replacing causal fusion with concatenation results in a decrement in DSC; removing multi-view transformations causes a drop. Together, these results validate our hypothesis that spatial causality is crucial for preserving semantic continuity in medical images.
@@ 5.3 Limitations and future directions
Despite the promising results, several challenges remain. First, integrating the Mamba macro-architecture into the SAMA module increases the parameter count, which highlights the need for more effective parameter optimization strategies. Second, extending the model to 3D medical image segmentation requires substantially higher memory, as it triples the number of scanning sequences in the Vision State Space module. To address this, future work can explore memory-efficient 3D processing strategies, such as patch-based or hybrid 2D–3D designs, along with model compression and pruning techniques to reduce resource demands without sacrificing accuracy. Designing lightweight yet expressive modules will be essential to balance efficiency and effectiveness, particularly for real-time 3D clinical applications. In future studies, we aim to refine the architecture, optimize parameter utilization, and investigate memory-efficient methods for scaling to large 3D datasets.
@@ 6 Conclusions
In this work, we presented SAMA-UNet, a novel architecture that integrates self-adaptive Mamba-like attention and causal multi-scale fusion for medical image segmentation. The proposed SAMA block and CR-MSM module enable the network to capture both fine local details and broad global context with improved efficiency. Extensive experiments on four benchmark datasets (BTCV, ACDC, EndoVis17, and ATLAS23) demonstrate that SAMA-UNet consistently outperforms CNN, Transformer, and Mamba-based models across CT, MRI, and endoscopic imaging. Future work will focus on developing lighter variants, exploring memory-efficient 3D adaptations, and enabling real-time deployment in clinical workflows.
@@
[…]
Acknowledgements This work was supported and funded by the Deanship of Scientific Research at Imam Mohammad Ibn Saud Islamic University (IMSIU) (grant number IMSIU-DDRSP2501)
[…]
Funding Statement This work was supported and funded by the Deanship of Scientific Research at Imam Mohammad Ibn Saud Islamic University (IMSIU) (grant number IMSIU-DDRSP2501).